### v1

@@ Abstract
keywords:
Causal Inference, Digital Twin, Industrial Control Systems, Cyber-Physical Security, Anomaly Detection, SWaT Testbed
@@ 1 Introduction
Industrial Control Systems (ICS) have changed from isolated proprietary networks to interconnected cyber-physical systems that combine operational technology with information technology infrastructure. This change started in the 1990s with the use of standard networking protocols and has grown faster with Industry 4.0 initiatives. Today, ICS need Internet connectivity for remote monitoring, predictive maintenance, and distributed control across facilities that are far apart [1, 2, 4, 5]. Critical infrastructure sectors such as water treatment, power generation, and manufacturing depend more on these connected systems, with the global ICS market expected to reach \$124 billion by 2025. The Secure Water Treatment (SWaT) testbed, developed by iTrust at Singapore University of Technology and Design, is an important example for cyber-physical security research. It provides a realistic scaled-down industrial environment with 51 sensors and actuators across six operational stages [3, 9, 13].
However, more connectivity and complexity create new cybersecurity vulnerabilities that traditional IT security cannot fully solve. The 2010 Stuxnet attack on Iranian nuclear facilities showed that advanced attackers can exploit both cyber weaknesses and physical process knowledge to cause real damage [6]. Recent events, like the 2021 Colonial Pipeline ransomware attack and the 2021 Oldsmar water treatment plant breach, show the ongoing threat to critical infrastructure. Current anomaly detection methods in ICS mostly use statistical correlation and machine learning techniques, which cannot separate real causal relationships from false associations [7, 8]. These correlation-based methods often have high false positive rates, cannot do root cause analysis, and are vulnerable to adversarial attacks that take advantage of missing causal understanding.
@@
In this research, we develop a comprehensive Causal Digital Twin (CDT) framework for cyber-physical security in medium-scale ICS. Our method combines causal inference theory with digital twin modeling to allow three types of causal reasoning: associational queries for pattern recognition, interventional queries to understand system reactions to defensive actions, and counterfactual queries for analyzing attacks retrospectively and planning prevention. We test our framework using the SWaT dataset, showing better performance than correlation-based methods over 41 cyber-physical attack scenarios, while also giving interpretable root cause analysis and actionable security recommendations.
The rest of this paper is organized as following. Section 2 gives a review of existing methods for ICS security and digital twin applications, and it shows the limitations of correlation-based methods. Section 3 defines the research problem and introduces the basic ideas of causal reasoning for cyber-physical security. Section 4 describes our new Causal Digital Twin framework, including causal graph discovery, estimation of structural causal model, and interventional reasoning algorithms. Section 5 shows the evaluation of the framework on the SWaT testbed compared with state-of-the-art methods. Finally, Section 6 finishes the paper with a summary of main contributions, discussion of practical implications, limitations of the work, and suggestions for future research.
@@ 2 Related Works
2.1 Digital Twins in Critical Infrastructure and ICS Security
@@
Digital twins have become an important tool for modeling and monitoring critical infrastructure systems, giving real-time virtual representations of physical assets and processes [16, 14, 15]. Xu et al. [2] made a detailed survey of digital twin applications in Industrial Internet of Things. They identified key technologies like data fusion, model synchronization, and real-time analytics for industrial automation. Their study shows that most current digital twin systems focus on operational optimization, not security, showing a clear gap in cybersecurity-focused digital twin frameworks. Qian et al. [17] proposed a full digital twin architecture for cyber-physical systems, highlighting the two-way data flow between physical and virtual parts. But their framework does not consider adversarial situations or causal reasoning, which are important for security use.
@@
Some recent work applied digital twins to security in specific areas. Coppolino et al. [19, 5] developed a cyber-resilient smart grid using digital twins and data spaces. They improved attack detection through continuous monitoring and anomaly detection. Their method improved threat detection accuracy but depends on predefined attack signatures and does not use causal reasoning to understand how attacks propagate. Erceylan et al. [20] suggested using digital twins for advanced threat modeling in cyber-physical systems, integrating security checks into the digital twin lifecycle. While this work gives important theoretical insights, it does not have empirical testing on real industrial data and does not include causal inference methods.
@@
Digital twin applications in industrial automation mostly focus on interoperability and data integration. Hildebrandt et al. [21] reviewed data integration methods for industrial digital twins, finding key needs for real-time synchronization, semantic interoperability, and scalable architectures. Platenius-Mohr et al. [22] showed interoperable digital twins in IIoT systems by transforming information models using Asset Administration Shell standards. However, these works focus on technical implementation and not on security or causal modeling.
@@
Some specialized applications in critical infrastructure show that digital twins can help cybersecurity. De Hoz Diego et al. [23] made an IoT digital twin for cybersecurity defense using runtime verification, achieving real-time threat detection by monitoring behavior continuously. Lampropoulos et al. [24] studied digital twins for critical infrastructure protection, emphasizing real-time situational awareness and automated responses. Sahoo et al. [25] proposed digital twin-enabled smart microgrids for full automation, improving resilience with predictive analytics and automated control. Even with these advances, none of these methods combine causal inference with digital twin modeling, so they cannot fully explain attack mechanisms or measure effectiveness of defensive actions.
2.2 Anomaly Detection Approaches in Industrial Control Systems
@@
Traditional anomaly detection in ICS has mostly used statistical and machine learning methods that detect deviations from normal operation. Kravchik and Shabtai [7] were among the first to use convolutional neural networks for ICS anomaly detection with the SWaT dataset. They achieved 86% detection accuracy by learning spatio-temporal patterns in sensor data. Their 1D convolution method worked better than recurrent neural networks, showing that deep learning can be effective for time-series anomaly detection in industrial systems. However, their approach only learns statistical patterns and does not understand physical processes or causal relationships. This makes it hard to separate normal operational changes from malicious attacks.
@@
Comparative studies show that anomaly detection performance varies across methods and datasets. Kim et al. [8, 38] evaluated five time-series anomaly detection models including InterFusion, RANSynCoder, GDN, LSTM-ED, and USAD on SWaT and HAI datasets. They found that InterFusion got the best F1-score on SWaT (90.7%), while RANSynCoder performed best on HAI (82.9%), showing that dataset characteristics affect performance. Their study also found that about 40% of training data is enough to reach near-optimal results, which is useful for practical deployment.
@@
Physics-informed methods try to combine data-driven learning with domain knowledge. Raman and Mathur [6] proposed PbNN, a hybrid physics-based neural network that uses ICS design knowledge to learn relationships among components. PbNN improved performance over purely data-driven methods by using physical constraints and process knowledge. Still, it relies on correlation-based learning and does not discover causal mechanisms, so it cannot fully explain attack causes or propagation paths.
@@
Recent work focused on explainability and feature importance to reduce the black-box problem in machine learning. Birihanu and Lendák [11] developed a correlation-based method using Latent Correlation Matrices and SHAP for root cause identification in ICS anomalies. Their method improved precision by 0.96% and gave interpretable explanations using SHAP values. Aslam et al. [12] combined Grey Wolf Optimizer with autoencoders to improve feature selection and model generalization, showing better accuracy, precision, recall, and F1-score on SWaT and WADI datasets. Even with these improvements, these approaches cannot separate causal factors from correlated confounders, which may mislead analysts about the real sources and mechanisms of attacks.
@@ 2.3 Causal Inference and Attribution Methods for ICS
Early works on causal methods for ICS security show that causal analysis can help to understand attacks and system weaknesses. Jadidi et al. [26] proposed ICS-CAD, a multi-step detection method that analyzes causal dependencies in ICS logs. Their method achieved 98% detection accuracy and could track how attacks propagate, showing the value of causal reasoning for multi-stage attacks. However, ICS-CAD considers only network-level causality and does not include causality of physical processes. This limits its use for cyber-physical attacks where sensors and actuators are manipulated while network behavior appears normal [31, 33, 34].
Causality-based approaches can also reduce false alarms. Koutroulis et al. [10] used causal models to capture interdependencies in water treatment systems and combined them with univariate models for single component behavior. Their method achieved the highest F1-score and zero false alarms on SWaT, performing better than correlation-based approaches. However, it depends strongly on domain knowledge to define causal links, which reduces generality and scalability for complex systems.
Studies on attribution show limits in ICS root cause analysis. Fung et al. [27] tested machine learning attribution methods on several datasets and found low performance because of timing problems, categorical actuators, and variability in attacks. Ensemble methods can help, but confounding variables and spurious correlations remain a problem.
Causal inference can also help to validate digital twins. Somers et al. [28] applied counterfactual analysis to confirm causal links in a robotic arm and detect faults. This shows potential for fault localization, but it focuses on model testing and does not consider realistic ICS security datasets.
@@ 2.4 Limitations and Research Gaps
Despite progress in digital twin technology and ICS security research, there are still important limitations that prevent effective causal reasoning for cyber-physical attacks. Current digital twin frameworks focus mostly on operational optimization and monitoring, and they do not include security-specific causal models. Existing anomaly detection methods can reach reasonable statistical performance, but they cannot distinguish true causal relations from spurious correlations. This limits root cause analysis and reduces effectiveness of defense strategies. Attribution and explainability methods are often unreliable for timing-sensitive and multi-modal cyber-physical attacks.
A key research gap is the lack of integration between causal inference theory and digital twins for comprehensive ICS security. No prior work provides a full framework for automated causal graph discovery from ICS data, structural causal model estimation for multi-stage industrial processes, and interventional or counterfactual reasoning for proactive security analysis. This gap prevents analysts from understanding the true causal mechanisms of attacks, reducing the effectiveness of incident response and limiting development of robust, causation-aware defenses.
Without causal reasoning, current methods cannot perform what-if analyses for defensive planning, analyze alternative attack scenarios, or reliably distinguish attack effects from normal operations. This leads to poor guidance for targeted mitigation, which is critical in industrial environments where wrong anomaly attribution can cause unnecessary shutdowns, delayed threat response, and ineffective security measures.
@@
CNN-based [7] Fast processing, effective pattern recognition No physical understanding, black-box nature None Statistical correlation only
Physics-informed (PbNN) [6] Incorporates domain knowledge, improved accuracy Still correlation-based, requires expert knowledge Limited Physics constraints only
Causality-inspired [10] Zero false alarms, domain-aware Manual causal specification, not generalizable Moderate Manual domain knowledge
SHAP-based [11] Feature importance, interpretable results Cannot distinguish causation from correlation High Attribution-based only
Digital Twin [13] Real-time monitoring, process understanding Pattern matching only, no causal reasoning Moderate Process simulation
ICS-CAD [26] Multi-step attack tracking, causal analysis Network-level only, limited to log analysis High Network causality only
@@ 3 Problem Statement
Industrial Control Systems (ICS) are becoming more vulnerable to advanced cyber-physical attacks that exploit both computational weaknesses and dependencies in physical processes. Traditional anomaly detection methods in critical infrastructure usually use correlation-based analysis. This type of analysis cannot separate true causal relations from spurious associations. Because of this limitation, the attack sources are often identified incorrectly and mitigation strategies are less effective [29].
The Secure Water Treatment (SWaT) testbed, created by iTrust at Singapore University of Technology and Design, provides a smaller but realistic water treatment environment. It has 51 sensors and actuators distributed over six operational stages [1]. The SWaT dataset includes 7 days of normal operation (around 500,000 samples) and 4 days of attack scenarios with 41 documented cyber-physical intrusions. It contains multivariate time-series data sampled at 1Hz and binary labels for normal or anomalous states.
3.1 Limitations of Correlation-Based Digital Twins
Existing digital twin frameworks for ICS have important limitations in causal reasoning. They cannot separate the three critical questions described in Pearl’s Causal Hierarchy [30, 35]:

### v2

@@ Abstract
Industrial Control Systems (ICS) in water distribution and treatment face cyber-physical attacks exploiting network and physical vulnerabilities. Current water system anomaly detection methods rely on correlations, yielding high false alarms and poor root cause analysis. We propose a Causal Digital Twin (CDT) framework for water infrastructures, combining causal inference with digital twin modeling. CDT supports association for pattern detection, intervention for system response, and counterfactual analysis for water attack prevention. Evaluated on water-related datasets SWaT, WADI, and HAI, CDT shows 90.8% compliance with physical constraints and structural Hamming distance 0.133 $\pm$ 0.02. F1-scores are $0.944\pm 0.014$ (SWaT), $0.902\pm 0.021$ (WADI), $0.923\pm 0.018$ (HAI, $p<0.0024$). CDT reduces false positives by 74%, achieves 78.4% root cause accuracy, and enables counterfactual defenses reducing attack success by 73.2%. Real-time performance at 3.2 ms latency ensures safe and interpretable operation for medium-scale water systems.
keywords:
@@
{graphicalabstract}
Graphical Overview [Uncaptioned image]
{highlights}
Novel CDT framework combines causal inference theory with digital twin modeling to enable associational, interventional, and counterfactual reasoning for cyber-physical security in ICS
Achieves statistically significant performance improvements with F1-scores of 0.944±0.014 (SWaT), 0.902±0.021 (WADI), and 0.923±0.018 (HAI), while reducing false positives by 74% compared to correlation-based methods
Automated causal structure discovery with 90.8% physical constraint compliance enables accurate root cause analysis achieving 78.4% Top-1 accuracy, improving 29.7% over traditional attribution methods
Counterfactual analysis capabilities provide proactive defense planning, identifying strategies that reduce attack success rates by 73.2% across 84 documented attack scenarios
Real-time deployment feasibility with 3.2ms inference latency and successful cross-dataset transfer (F1-scores 0.847-0.889), demonstrating industrial scalability for medium-scale control systems.
@@ 1 Introduction
The ICS have changed from isolated proprietary networks to interconnected cyber-physical systems that combine operational technology with information technology infrastructure. This change started in the 1990s with the use of standard networking protocols and has grown faster with Industry 4.0 initiatives. Today, ICS need Internet connectivity for remote monitoring, predictive maintenance, and distributed control across facilities that are far apart Mathur & Tippenhauer (2016); Xu et al. (2023); Abdoli & Ye (2023); Jiang et al. (2024). Critical infrastructure sectors such as water treatment, power generation, and manufacturing depend more on these connected systems, with the global ICS market expected to reach \$124 billion by 2025. The Secure Water Treatment (SWaT) testbed, developed by iTrust at Singapore University of Technology and Design, is an important example for cyber-physical security research. It provides a realistic scaled-down industrial environment with 51 sensors and actuators across six operational stages Goh et al. (2017); Jeremiah et al. (2024); Lucchese et al. (2024).
However, more connectivity and complexity create new cybersecurity vulnerabilities that traditional IT security cannot fully solve. The 2010 Stuxnet attack on Iranian nuclear facilities showed that advanced attackers can exploit both cyber weaknesses and physical process knowledge to cause real damage Raman & Mathur (2022). Recent events, like the 2021 Colonial Pipeline ransomware attack and the 2021 Oldsmar water treatment plant breach, show the ongoing threat to critical infrastructure. Current anomaly detection methods in ICS mostly use statistical correlation and machine learning techniques, which cannot separate real causal relationships from false associations Kravchik & Shabtai (2018); Kim et al. (2023). These correlation-based methods often have high false positive rates, cannot do root cause analysis, and are vulnerable to adversarial attacks that take advantage of missing causal understanding.
@@
In this research, we develop a comprehensive CDT framework for cyber-physical security in medium-scale ICS. Our method combines causal inference theory with DT modeling to allow three types of causal reasoning: associational queries for pattern recognition, interventional queries to understand system reactions to defensive actions, and counterfactual queries for analyzing attacks retrospectively and planning prevention. We test our framework using the SWaT dataset, showing better performance than correlation-based methods over 41 cyber-physical attack scenarios, while also giving interpretable root cause analysis and actionable security recommendations.
The rest of this paper is organized as following. Section 2 gives a review of existing methods for ICS security and DT applications, and it shows the limitations of correlation-based methods. Section 3 defines the research problem and introduces the basic ideas of causal reasoning for cyber-physical security. Section 4 describes our newCDT framework, including causal graph discovery, estimation of structural causal model, and interventional reasoning algorithms. Section 5 shows the evaluation of the framework on the SWaT testbed compared with state-of-the-art methods. Finally, Section 6 finishes the paper with a summary of main contributions, discussion of practical implications, limitations of the work, and suggestions for future research.
@@ 2 Related Works
2.1 DTs in Critical Infrastructure and ICS Security
@@
DTs have become an important tool for modeling and monitoring critical infrastructure systems, giving real-time virtual representations of physical assets and processes Homaei et al. (2022, 2024b, 2024a). Xu et al. Xu et al. (2023) made a detailed survey of DT applications in Industrial Internet of Things (IoT). They identified key technologies like data fusion, model synchronization, and real-time analytics for industrial automation. Their study shows that most current DT systems focus on operational optimization, not security, showing a clear gap in cybersecurity-focused DT frameworks. Qian et al. Qian et al. (2022) proposed a full DT architecture for cyber-physical systems, highlighting the two-way data flow between physical and virtual parts. But their framework does not consider adversarial situations or causal reasoning, which are important for security use.
@@
Some recent work applied DTs to security in specific areas. Coppolino et al. Coppolino et al. (2023); Jiang et al. (2024) developed a cyber-resilient smart grid using DTs and data spaces. They improved attack detection through continuous monitoring and anomaly detection. Their method improved threat detection accuracy but depends on predefined attack signatures and does not use causal reasoning to understand how attacks propagate. Erceylan et al. Erceylan et al. (2025) suggested using DTs for advanced threat modeling in cyber-physical systems, integrating security checks into the DT lifecycle. While this work gives important theoretical insights, it does not have empirical testing on real industrial data and does not include causal inference methods.
@@
DT applications in industrial automation mostly focus on interoperability and data integration. Hildebrandt et al. Hildebrandt et al. (2024) reviewed data integration methods for industrial DTs, finding key needs for real-time synchronization, semantic interoperability, and scalable architectures. Platenius-Mohr et al. Platenius-Mohr et al. (2019) showed interoperable DTs in IIoT systems by transforming information models using Asset Administration Shell standards. However, these works focus on technical implementation and not on security or causal modeling.
@@
Some specialized applications in critical infrastructure show that DTs can help cybersecurity. De Hoz Diego et al. de Hoz Diego et al. (2022) made an IoT DT for cybersecurity defense using runtime verification, achieving real-time threat detection by monitoring behavior continuously. Lampropoulos et al. Lampropoulos et al. (2024) studied DTs for critical infrastructure protection, emphasizing real-time situational awareness and automated responses. Sahoo et al. Sahoo et al. (2025) proposed DT-enabled smart microgrids for full automation, improving resilience with predictive analytics and automated control. Even with these advances, none of these methods combine causal inference with DT modeling, so they cannot fully explain attack mechanisms or measure effectiveness of defensive actions.
2.2 Anomaly Detection Approaches in ICS
@@
Traditional anomaly detection in ICS has mostly used statistical and machine learning methods that detect deviations from normal operation. Kravchik and Shabtai Kravchik & Shabtai (2018) were among the first to use convolutional neural networks for ICS anomaly detection with the SWaT dataset. They achieved 86% detection accuracy by learning spatio-temporal patterns in sensor data. Their 1D convolution method worked better than recurrent neural networks, showing that deep learning can be effective for time-series anomaly detection in industrial systems. However, their approach only learns statistical patterns and does not understand physical processes or causal relationships. This makes it hard to separate normal operational changes from malicious attacks.
@@
Comparative studies show that anomaly detection performance varies across methods and datasets. Kim et al. Kim et al. (2023); Mogensen et al. (2024) evaluated five time-series anomaly detection models including InterFusion, RANSynCoder, GDN, LSTM-ED, and USAD on SWaT and HAI datasets. They found that InterFusion got the best F1-score on SWaT (90.7%), while RANSynCoder performed best on HAI (82.9%), showing that dataset characteristics affect performance. Their study also found that about 40% of training data is enough to reach near-optimal results, which is useful for practical deployment.
@@
Physics-informed methods try to combine data-driven learning with domain knowledge. Raman and Mathur Raman & Mathur (2022) proposed Physics-based Neural Network (PbNN), a hybrid physics-based neural network that uses ICS design knowledge to learn relationships among components. PbNN improved performance over purely data-driven methods by using physical constraints and process knowledge. Still, it relies on correlation-based learning and does not discover causal mechanisms, so it cannot fully explain attack causes or propagation paths.
@@
Recent work focused on explainability and feature importance to reduce the black-box problem in machine learning. Birihanu and Lendák Birihanu & Lendák (2025) developed a correlation-based method using Latent Correlation Matrices and SHAP for root cause identification in ICS anomalies. Their method improved precision by 0.96% and gave interpretable explanations using SHAP values. Aslam et al. Aslam et al. (2025) combined Grey Wolf Optimizer with autoencoders to improve feature selection and model generalization, showing better accuracy, precision, recall, and F1-score on SWaT and WADI datasets. Even with these improvements, these approaches cannot separate causal factors from correlated confounders, which may mislead analysts about the real sources and mechanisms of attacks.
@@ 2.3 Causal Inference and Attribution Methods for ICS
Early works on causal methods for ICS security show that causal analysis can help to understand attacks and system weaknesses. Jadidi et al. Jadidi et al. (2022) proposed ICS-CAD, a multi-step detection method that analyzes causal dependencies in ICS logs. Their method achieved 98% detection accuracy and could track how attacks propagate, showing the value of causal reasoning for multi-stage attacks. However, ICS-CAD considers only network-level causality and does not include causality of physical processes. This limits its use for cyber-physical attacks where sensors and actuators are manipulated while network behavior appears normal Peters et al. (2017); Imbens & Rubin (2015); Hernán & Robins (2024).
Causality-based approaches can also reduce false alarms. Koutroulis et al. Koutroulis et al. (2022) used causal models to capture interdependencies in water treatment systems and combined them with univariate models for single component behavior. Their method achieved the highest F1-score and zero false alarms on SWaT, performing better than correlation-based approaches. However, it depends strongly on domain knowledge to define causal links, which reduces generality and scalability for complex systems.
Studies on attribution show limits in ICS root cause analysis. Fung et al. Fung et al. (2024) tested machine learning attribution methods on several datasets and found low performance because of timing problems, categorical actuators, and variability in attacks. Ensemble methods can help, but confounding variables and spurious correlations remain a problem.
Causal inference can also help to validate DTs. Somers et al. Somers et al. (2022) applied counterfactual analysis to confirm causal links in a robotic arm and detect faults. This shows potential for fault localization, but it focuses on model testing and does not consider realistic ICS security datasets.
@@ 2.4 Limitations and Research Gaps
Although DT technology and ICS security research made progress, there are still big limits for causal reasoning on cyber-physical attacks. Most DT frameworks focus on optimization and monitoring, but they do not include security-specific causal models. Current anomaly detection shows good results, but it cannot separate real causal relations from false correlations, which reduces root cause analysis and defense. Attribution and explainability are also not reliable for attacks that are fast and multi-modal.
The main research gap is missing integration of causal inference with DTs for ICS security. No work until now gives a full framework for causal graph discovery from ICS data, estimation of causal models for multi-stage processes, and interventional or counterfactual reasoning for proactive analysis. Because of this gap, analysts cannot fully understand attack mechanisms, which reduces effective incident response and blocks the development of causation-aware defenses (Table 1). Without causal reasoning, current methods cannot do what-if analysis, test alternative attack scenarios, or clearly separate attack effects from normal operation. This gives weak guidance for mitigation, which is dangerous in industrial systems where wrong attribution can cause shutdowns, delays in response, and weak security actions.
@@
CNN-based
Kravchik and Shabtai (2018) Fast processing, effective pattern recognition No physical understanding, black-box nature None Statistical correlation only
Physics-informed
Raman (2022) Incorporates domain knowledge, improved accuracy Still correlation-based, requires expert knowledge Limited Physics constraints only
Causality-inspired
Koutroulis et al. (2022) Zero false alarms, domain-aware Manual causal specification, not generalizable Moderate Manual domain knowledge
SHAP-based
Birihanu et al. (2025) Feature importance, interpretable results Cannot distinguish causation from correlation High Attribution-based only
Digital Twin
Lucchese et al. (2024) Real-time monitoring, process understanding Pattern matching only, no causal reasoning Moderate Process simulation
ICS-CAD
Jadidi et al. (2022) Multi-step attack tracking, causal analysis Network-level only, limited to log analysis High Network causality only
@@ 3 Problem Statement
The ICS are becoming more vulnerable to advanced cyber-physical attacks that exploit both computational weaknesses and dependencies in physical processes. Traditional anomaly detection methods in critical infrastructure usually use correlation-based analysis. This type of analysis cannot separate true causal relations from spurious associations. Because of this limitation, the attack sources are often identified incorrectly and mitigation strategies are less effective Pearl (2009).
The SWaT testbed, created by iTrust at Singapore University of Technology and Design, provides a smaller but realistic water treatment environment. It has 51 sensors and actuators distributed over six operational stages Mathur & Tippenhauer (2016). The SWaT dataset includes 7 days of normal operation (around 500,000 samples) and 4 days of attack scenarios with 41 documented cyber-physical intrusions. It contains multivariate time-series data sampled at 1Hz and binary labels for normal or anomalous states.
3.1 Limitations of Correlation-Based DTs
Existing DT frameworks for ICS have important limitations in causal reasoning. They cannot separate the three critical questions described in Pearl’s Causal Hierarchy Pearl & Mackenzie (2019); Glymour et al. (2019):